# ToDA: Target-oriented Diffusion Attacker against Recommendation System

### v1

@@ 3.1. Latent Diffusion Attacker
DMs are employed as the foundational attacker model for user profile generation in our approach. In a typical DMs’ process, noise is gradually added at each forward step, while the reverse operation attempts to reconstruct the state from the preceding step. These processes are guided by the Markov assumption, implying that each state is exclusively reliant on its immediate predecessor. Consequently, these two processes can be mathematically represented as follows:
@@
(1) $q(\mathbf{Y}_{u;1:S}|\mathbf{Y}_{u;0})=\prod_{s=1}^{S}q(\mathbf{Y}_{u;s}|\mathbf{Y}_{u;s-1}),$
@@
(2) $p(\mathbf{Y}_{u;0:S})=p(\mathbf{Y}_{u;0})\prod_{s=1}^{S}p(\mathbf{Y}_{u;s-1}|\mathbf{Y}_{u;s}),$
@@
where $p(\mathbf{Y}_{u;0:S})$ denotes the probability distribution of the sequence of user profiles during the reverse process. The term $p(\mathbf{Y}_{u;0})$ stands for the probability of the initial profile (i.e., $\mathbf{Y}_{u;0}=\mathbf{Y}_{u}$), while the product term represents the conditional probability of each preceding profile $\mathbf{Y}_{u;s-1}$ given its successor $\mathbf{Y}_{u;s}$, iteratively applied over all $S$ steps.
Figure 2. The overall framework of ToDA. It is featured by the latent diffusion attacker with target-oriented approximator $\boldsymbol{\epsilon}_{\theta}$. The approximator $\boldsymbol{\epsilon}_{\theta}$ denoises the previous state of user feature $\mathbf{e}_{u;s}$, bringing in the global information to the local view diffusion model.
@@
To facilitate the malicious profile generation and diminish the computational overhead during training, we adopt principles akin to perceptual image compression in Computer Vision (Rombach et al., 2022), exploiting the benefits of high-dimensional spaces. More precisely, given a user profile $\mathbf{Y}_{u}\in\{0,1\}^{m}$, the encoder $\mathcal{E}$ encodes $\mathbf{Y}_{u}$ into a latent representation $\mathbf{e}_{u}=\mathcal{E}(\mathbf{Y}_{u})\in\mathbb{R}^{d}$ and the decoder $\mathcal{D}$ reconstructs the user profile $\mathbf{Y}^{a}_{u}=\mathcal{D}(\mathbf{e}_{u})$. Thereby we construct the latent diffusion attacker inheriting the Equation 1 and 2 as $q(\mathbf{e}_{u;1:S}|\mathbf{e}_{u;0})$ and $p(\mathbf{e}_{u;0:S})$, respectively.
To optimize the parameters of the latent diffusion models, we employ multinomial likelihood as the basis for these elements (Liang et al., 2018), giving us the following expression for $\mathcal{D}$ and $\mathcal{E}$ as:
(3) $\mathcal{L}_{r}=-\mathbb{E}_{i\in\mathcal{I}}\mathbf{Y}_{u,i}\log{\mathcal{D}(\mathbf{e}_{u})},$
As depicted in Figure 2, the user profiles are transformed into a dense feature, slightly diffused with noises step by step. In reversion, Target-oriented Approximator serves to recover the original features and incorporate target items to guide the generation procedure.
In this case, we define our diffusion objective, $\mathcal{L}_{d}$, as follows:
(4) $\mathcal{L}_{d}$ $=\mathbb{E}_{s\sim[1,S],\mathbf{e}_{u;0},\boldsymbol{\epsilon}_{s}}\Big{[}\|\boldsymbol{\epsilon}_{s}-\boldsymbol{\epsilon}_{\theta}(\mathbf{e}_{u;s},s)\|^{2}\Big{]}$
$=\mathbb{E}_{s\sim[1,S],\mathbf{e}_{u;0},\boldsymbol{\epsilon}_{s}}\Big{[}\|\boldsymbol{\epsilon}_{s}-\boldsymbol{\epsilon}_{\theta}(\sqrt{\bar{\alpha}_{s}}\mathbf{e}_{u;0}+\sqrt{1-\bar{\alpha}_{s}}\boldsymbol{\epsilon}_{s},s)\|^{2}\Big{]},$
@@
where we calculate the Mean Squared Error (MSE) between the noise $\boldsymbol{\epsilon}_{s}$ and the estimated one approximated by $\boldsymbol{\epsilon}_{\theta}$. We exploit a reparameterization trick that allows for a closed-form solution for the forward procedure (Ho et al., 2020). This means that we can directly compute the variable at a future step $s$, $\mathbf{e}_{u;s}$, from the initial variable $\mathbf{e}_{u;0}$ and the variable $\boldsymbol{\epsilon}_{s}$. This is derived from the formula as
(5) $q(\mathbf{e}_{u;s}|\mathbf{e}_{u;s-1})=\mathcal{N}(\mathbf{e}_{u;s};\sqrt{1-\beta_{s}}\mathbf{e}_{u;s-1},\beta_{s}\mathbf{I}),$
where $\{\beta_{s}\}^{S}_{s=1}$ is the variance schedule to control the noise scale. Let $\alpha_{s}=1-\beta_{s}$ and $\bar{\alpha}_{s}=\prod_{i=1}^{s}\alpha_{i}$, we have
(6) $q(\mathbf{e}_{u;s}|\mathbf{e}_{u;0})=\mathcal{N}(\mathbf{e}_{u;s};\sqrt{\bar{\alpha}_{t}}\mathbf{e}_{u;0},(1-\bar{\alpha}_{s})\mathbf{I}),$
which makes it effectively to produce $\mathbf{e}_{u;s}$ at any steps.
3.2. Target-oriented Approximator
Figure 3. A denoise step of ToDA, depicting the target-oriented approximator $\boldsymbol{\epsilon}_{\theta}$. The features of target items are incorporated through a GCN Encoder $\mathcal{E}_{\mathcal{G}}$, before using the cross-attention to control the denoising orientation. The internal autoencoder is leveraged to compress the original feature into a bottleneck $\mathbf{b}_{u;s}$.
@@ 3.2. Target-oriented Approximator
As aforementioned, a conventional attacker generates the user profile without consideration of the correlations between target items and auxiliary items, leading to a dispersion of attack targets and an unintentional promotion. To endow the diffusion attacker with the target-oriented ability, we tailor the estimator $\epsilon_{\theta}$ with a global horizon view and leverage the cross attention to transfer the target information into the latent user feature.
@@
To obtain the global view item feature, we commerce with using the encoded user features $\mathbf{e}_{u}$ to initialize the item embeddings as:
(7) $\mathbf{e}_{i}^{0}=\frac{1}{|\mathcal{N}_{i}|}\sum_{u\in\mathcal{N}_{i}}\mathbf{e}_{u},$
where $\mathcal{N}_{i}$ is the indices of users that are interacted with item $i$. By aggregating the interacted user features, we maintain a consistent space alignment with the latent diffusion attacker, facilitating the space alignment and following transferring.
@@
Inspired by the high-order information capturing, we apply a GNN model to learn the item features. The GNN model is adept at processing graph-structured data, thereby enabling the extraction of informative representations for nodes in a graph. This property is particularly beneficial in our case, where interactions among users and items can be naturally modeled as a graph. Moreover, this global view extends the generation of a broader horizon outwards the on-processing user feature itself, compensating the narrow focus of DMs. Formally, we learn the target item features as:
(8) $\mathbf{e}_{t}=\mathcal{E}_{\mathcal{G}}(t),\mathcal{G}=(\mathbf{e}_{\mathcal{U}},\mathbf{e}_{\mathcal{I}}^{0},\mathbf{Y}),$
@@
where $\mathcal{E}_{\mathcal{G}}$ is the GNN encoder whose input is the user-item bipartite graph $\mathcal{G}$ and the index of the target item $t$. $\mathbf{e}_{\mathcal{I}}^{0}$ is the matrix consisting of all item embeddings generated by Equation  7.
@@
In the context of DMs, an approximator is employed to estimate noise and progressively refines the input $\mathbf{e}_{u;S}$. While this method is exquisite, it overlooks the significance of target items. Its primary focus remains ensuring that shilling attacks go unnoticed. To rectify this limitation and bolster the efficacy of the attacks, we enhance the original approximator, denoted as $\epsilon_{\theta}$, to focus on the target items. This modified approximate takes into consideration the user’s latent feature $\mathbf{e}_{u;s}$ at the preceding step $s$ and the target item feature $\mathbf{e}_{t}$, then producing the next user latent feature $\mathbf{e}_{u;s-1}$. For the purpose of efficiently handling the information related to target items, we employ AE models, which allows us to compress the original feature with $\mathcal{E}^{\prime}$ into a bottleneck $\mathbf{b}_{u;s}$, and then to reconstruct it using $\mathcal{D}^{\prime}$. To pre-process $\mathbf{e}_{t}$, we introduce a transformation $\tau$ that projects $\mathbf{e}_{t}$ to $\tau(\mathbf{e}_{t})$ aligned with the $\mathbf{b}_{u;s}$, which is then carried out through a cross-attention layer implementing $Attn(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{softmax}(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{d}})\cdot\mathbf{V}$, with
(9) $\mathbf{Q}=\mathbf{W}_{Q}\cdot\mathbf{b}_{u;s},\ \mathbf{K}=\mathbf{W}_{K}\cdot\tau(\mathbf{e}_{t}),\ \mathbf{V}=\mathbf{W}_{V}\cdot\tau(\mathbf{e}_{t}),$
where $\mathbf{W}_{Q},\mathbf{W}_{K},\mathbf{W}_{V}\in\mathbb{R}^{d^{\prime}\times d}$ are learnable matrices. A visual representation of the denoising step $s$ is provided in Figure 3. By performing $S$ steps of this target-oriented noise approximation, we obtain the fake user profile $\mathbf{Y}^{a}_{u}$ at the end which reaps the benefits of both imperceptibility and attack-effectiveness.
3.3. Optimization

### v3

@@ III-B Latent Diffusion Attacker
DMs are employed as the foundational attacker model for user profile generation in our approach. In a typical DMs’ process, noise is gradually added at each forward step, while the reverse operation attempts to reconstruct the state from the preceding step. These processes are guided by the Markov assumption, implying that each state is exclusively reliant on its immediate predecessor. Consequently, these two processes can be formally represented as follows:
@@
$q(\mathbf{Y}_{u;1:S}|\mathbf{Y}_{u;0})=\prod_{s=1}^{S}q(\mathbf{Y}_{u;s}|\mathbf{Y}_{u;s-1}),$ (3)
@@
$p(\mathbf{Y}_{u;0:S})=p(\mathbf{Y}_{u;0})\prod_{s=1}^{S}p(\mathbf{Y}_{u;s-1}|\mathbf{Y}_{u;s}),$ (4)
@@
where $p(\mathbf{Y}_{u;0:S})$ denotes the probability distribution of the sequence of user profiles during the reverse process. The term $p(\mathbf{Y}_{u;0})$ stands for the probability of the initial profile (i.e., $\mathbf{Y}_{u;0}=\mathbf{Y}_{u}$), while the product term represents the conditional probability of each preceding profile $\mathbf{Y}_{u;s-1}$ given its successor $\mathbf{Y}_{u;s}$, iteratively applied over all $S$ steps. Following these procedure, The conventional DMs achieve reconstruction of the user profiles in a discrete state, to ensure the imperceptibility. However, such discrete state has a large process space which aligns with the size of items, thus increasing the computational burden and further hindering the flexibility of conditional generation for DMs.
@@
To facilitate the malicious profile generation and diminish the computational overhead during training, we adopt principles akin to perceptual image compression in computer vision [41], exploiting the benefits of high-dimensional spaces. More precisely, given a user profile $\mathbf{Y}_{u}\in\{0,1\}^{m}$, the encoder $\mathcal{E}$ encodes $\mathbf{Y}_{u}$ into a latent representation $\mathbf{e}_{u}=\mathcal{E}(\mathbf{Y}_{u})\in\mathbb{R}^{d}$ and the decoder $\mathcal{D}$ reconstructs the user profile $\mathbf{Y}^{a}_{u}=\mathcal{D}(\mathbf{e}_{u})$. Wherein, both the encoder $\mathcal{E}$ and decoder $\mathcal{D}$ are implemented by the multilayer perceptron (MLP), following the architecture of MultiDAE [42]. And we calculate the multinomial likelihood to optimize the parameters of $\mathcal{D}$ and $\mathcal{E}$, formally,
$\mathcal{L}_{r}=-\mathbb{E}_{i\in\mathcal{I}}\mathbf{Y}_{u,i}\log{\mathcal{D}(\mathbf{e}_{u})}.$ (5)
After the training via $\mathcal{L}_{r}$, we obtain the paired autoencoder that is capable to transform discrete user profiles into continual user features. Thereby we construct the latent diffusion attacker inheriting the Equation 3 and 4 as $q(\mathbf{e}_{u;1:S}|\mathbf{e}_{u;0})$ and $p(\mathbf{e}_{u;0:S})$, respectively.
For the forward process, we add noise to the user feature step by step following:
$q(\mathbf{e}_{u;s}|\mathbf{e}_{u;s-1})=\mathcal{N}(\mathbf{e}_{u;s};\sqrt{1-\beta_{s}}\mathbf{e}_{u;s-1},\beta_{s}\mathbf{I}),$ (6)
@@
where $\{\beta_{s}\}^{S}_{s=1}$ is the variance schedule to control the noise scale. And we exploit a reparameterization trick  [43] to formulate the distribution after adding the gaussian noise. Such trick also allows for a closed-form solution for the forward procedure, which means that we can directly compute the user features at a future step $s$, $\mathbf{e}_{u;s}$, from the initial state $\mathbf{e}_{u;0}$:
$q(\mathbf{e}_{u;s}|\mathbf{e}_{u;0})=\mathcal{N}(\mathbf{e}_{u;s};\sqrt{\bar{\alpha}_{t}}\mathbf{e}_{u;0},(1-\bar{\alpha}_{s})\mathbf{I}),$ (7)
where $\bar{\alpha}_{s}=\prod_{i=1}^{s}(1-\beta_{i})$. For the reverse process, we aim to iteratively denoise $\mathbf{e}_{u;s}$ for $s$ steps to obtain the initial user feature $\mathbf{e}_{u;0}$:
$p(\mathbf{e}_{u;s-1}|\mathbf{e}_{u;s})=\mathcal{N}(\mathbf{e}_{u;s};\boldsymbol{\epsilon}_{\theta}(\mathbf{e}_{u;s},s,\mathbf{e}_{t}),\Sigma_{\theta}(\mathbf{e}_{u;s},s)),$ (8)
where $\Sigma_{\theta}(\mathbf{e}_{u;s},s)=\frac{1-\bar{\alpha}_{s-1}}{1-\bar{\alpha}_{s}}\beta_{s}\mathbf{I}$ denotes the variance and the target-oriented approximator $\boldsymbol{\epsilon}_{\theta}(\cdot)$ takes the target item feature $\mathbf{e}_{t}$ as the conditional signal to guide the reversion procedure. To optimize the latent diffusion model, DMs compel the posterior distribution closer to the prior distribution via KL divergence:
$\mathcal{L}_{d}$ $=\mathbb{E}_{s\sim[1,S],\mathbf{e}_{u;0},\boldsymbol{\epsilon}_{s}}\text{KL}(q(\mathbf{e}_{u;s-1}|\mathbf{e}_{u;0})||p(\mathbf{e}_{u;s-1}|\mathbf{e}_{u;s}))$ (9)
Thanks to the DDPM framework [43], the above objective can be simplified to a Mean Squared Error (MSE) between the noise $\boldsymbol{\epsilon}_{s}$ and the estimated one approximated by $\boldsymbol{\epsilon}_{\theta}$ as follows:
$\mathcal{L}_{d}$ $=\mathbb{E}_{s\sim[1,S],\mathbf{e}_{u;0},\boldsymbol{\epsilon}_{s}}\Big{[}\|\boldsymbol{\epsilon}_{s}-\boldsymbol{\epsilon}_{\theta}(\mathbf{e}_{u;s},s,\mathbf{e}_{t})\|^{2}\Big{]}$ (10)
$=\mathbb{E}_{s\sim[1,S],\mathbf{e}_{u;0},\boldsymbol{\epsilon}_{s}}\Big{[}\|\boldsymbol{\epsilon}_{s}-\boldsymbol{\epsilon}_{\theta}(\sqrt{\bar{\alpha}_{s}}\mathbf{e}_{u;0}+\sqrt{1-\bar{\alpha}_{s}}\boldsymbol{\epsilon}_{s},s,\mathbf{e}_{t})\|^{2}\Big{]}.$
III-C Target-oriented Approximator
@@
The latent diffusion model achieves efficient generation of user profiles, while lacking the malicious intent. To endow the diffusion attacker with the target-oriented ability, we tailor the estimator $\epsilon_{\theta}$ with a global horizon view to compensate the narrow focus of DMs and leverage the cross attention to transfer the target information into the latent user feature.
@@
To obtain the global view item feature, we commerce with using the encoded user profile features $\mathbf{e}_{u}$ to initialize the item embeddings as:
$\mathbf{e}_{i}=\frac{1}{|\mathcal{N}_{i}|}\sum_{u\in\mathcal{N}_{i}}\mathbf{e}_{u},$ (11)
@@
where $\mathcal{N}_{i}$ is the indices of users that are interacted with item $i$. By aggregating the interacted user features, we maintain a consistent space alignment with the latent diffusion attacker, facilitating the space alignment and following transferring. Inspired by the high-order information capturing, we apply a GNN model to learn the item features from the message-passing mechanism. The GNN model is adept at processing graph-structured data, thereby enabling the extraction of informative representations for nodes, where interactions among users and items can be naturally modeled as a graph, denoted as $\mathcal{G}=(\{\mathcal{U},\mathcal{I}\},\mathbf{Y})$. $\{\mathcal{U},\mathcal{I}\}$ denotes the nodes while $\mathbf{Y}$ represents the edges of graph $\mathcal{G}$. Obviously, the integration of graphs extends the generation of a broader horizon outwards the on-processing user feature itself, compensating the narrow focus of DMs. Formally, we learn the target item features as:
$\mathbf{e}_{t}=\mathcal{E}_{\mathcal{G}}(t),\quad\mathcal{G}=(\mathbf{e}_{\mathcal{U}},\mathbf{e}_{\mathcal{I}},\mathbf{Y}),$ (12)
@@
where $\mathcal{E}_{\mathcal{G}}$ is the GNN encoder, which is equipped with the user-item bipartite graph $\mathcal{G}$, and takes the index of the target item $t$ as input. $\mathbf{e}_{\mathcal{I}}$ is the matrix consisting of all item embeddings generated by Equation  11. Specifically, we remove the parameters of conventional GCN to achieve efficient feature aggregation via normalized symmetric adjacency matrix: $\hat{\mathbf{A}}=\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}$ [3], therefore seamlessly obtaining the target item features from the global structure of user-item interactions.
@@
In the context of DMs, an approximator is employed to estimate noise and progressively refines the input $\mathbf{e}_{u;S}$. While this method is exquisite, it overlooks the significance of target items. Its primary focus remains ensuring that shilling attacks go unnoticed. To rectify this limitation and bolster the efficacy of the attacks, we enhance the original approximator, denoted as $\epsilon_{\theta}$, to focus on the target items, reaping the benefits of both efficacy and efficiency. This modified approximate takes into consideration the user’s latent feature $\mathbf{\hat{e}}_{u;s}$ at the preceding step $s$ and the target item feature $\mathbf{e}_{t}$, then producing the next user latent feature $\mathbf{\hat{e}}_{u;s-1}$. For the purpose of efficiently handling the information related to target items, we employ AE models, which allows us to compress the original feature with $\mathcal{E}^{\prime}$ into a latent bottleneck $\mathbf{b}_{u;s}$, and then to reconstruct it using $\mathcal{D}^{\prime}$. To pre-process $\mathbf{e}_{t}$, we introduce a transformation $\tau$ that projects $\mathbf{e}_{t}$ to $\tau(\mathbf{e}_{t})$ aligned with the $\mathbf{b}_{u;s}$, which is then carried out through a cross-attention layer implementing
$Attn(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{softmax}(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{d}})\cdot\mathbf{V},$ (13)
$\mathbf{Q}=$ $\mathbf{W}_{Q}\cdot\mathbf{b}_{u;s},\ \mathbf{K}=\mathbf{W}_{K}\cdot\tau(\mathbf{e}_{t}),\ \mathbf{V}=\mathbf{W}_{V}\cdot\tau(\mathbf{e}_{t}),$
where $\mathbf{W}_{Q},\mathbf{W}_{K},\mathbf{W}_{V}\in\mathbb{R}^{d^{\prime}\times d}$ are learnable matrices. Note that, our method is also flexible to incorporate multiple target items, which can be interpreted in a matrix format: $\mathbf{Q}\in\mathbb{R}^{|\mathcal{U}^{\prime}|\times d^{\prime}}$, and $\mathbf{K},\mathbf{V}\in\mathbb{R}^{|\mathcal{T}|\times d^{\prime}}$, where $|\mathcal{U}^{\prime}|$ represents a batch of users and $|\mathcal{T}|$ denotes the number of target items. Therefore, $\mathbf{Q}\times\mathbf{K}^{T}\in\mathbb{R}^{|\mathcal{U}^{\prime}|\times|\mathcal{T}|}$, leading to the final representation $Attn(\mathbf{Q},\mathbf{K},\mathbf{V})\in\mathbb{R}^{|\mathcal{U}^{\prime}|\times d^{\prime}}$, which is guided by multiple target items from $\mathcal{T}$. A visual representation of the denoising step $s$ is provided in Figure 2. By performing $S$ steps of this target-oriented noise approximation, we obtain the fake user profile $\mathbf{Y}^{a}_{u}$ at the end which reaps the benefits of both imperceptibility and attack-effectiveness. Notably, conventional DMs typically employ complicated and heavy modules such as U-Net [44] as approximators, constrained by significant computational demands. In contrast, our method is notably more efficient, balancing lightweight design with effectiveness. Further analysis on this is detailed in Section III-E.
III-D Optimization